The pulse rates of 200 college men are normally distributed with a mean of 72 and standard deviation 6. Using the empirical rule, find
(a) the percentage of college men whose pulse rates are between 66 and 84. Do not write the percent sign.

(b) the percentage of college men whose pulse rates are at most 60. Do not write the percent sign.

(c) the number of college men whose pulse rates are at least 78. Do not give answer as a percent.
Answer #1

x ~110,0) N Empirical fule Mto = 68% 4720 95% free Чt 3r 99.7% 13434 y 73.0 is 3-2 2.35 2 3 2.35b P(66<x<84) 66-72 < Zc 6 :P(66 84772) =P(-1<2<2) 68 +13 = b P(X360) EP (25 60-72 ) 6 =P (24-2) 50 - 34-13.5 -2.50 pl x7,78)